# methodsReferences for Methods

### v1

@@ Introduction
We now turn to a set of realistic models of nuclear binding energy. For this study, we have chosen 15 realistic computational models that represent a few classes of theoretical frameworks that see broad use. The models are specified in Table 2, with their respective RMSE for each of the three datasets. As the model orthogonalization and mixing strategies only require precomputed data across a range of nuclei, we pull forecasts directly from published datasets. The specific mass models chosen are those from the MassExplorer database [32] and HFB-24[33] and FRDM-12 [34] mass tables.
@@
The domains of the experimental datasets used are shown in Fig. 3(a). As in the case study I, we divide the 629 data points into training, validation, and testing sets. We perform the SVD on the forecasts produced by the set of 15 models restricted to the training set and show their projections $\nu_{1}$ and $\nu_{2}$ on the first two principal components in Fig. 3(b). The nearly-exponential decay of the singular values (6) is shown in the inset of panel (c). Note that the SVD is performed after the centering procedure, thus the principal components in panel (b) should be interpreted as the variability about the average in the space of predictions. While it is tempting to draw conclusions on model similarity from these projections, it can only be said that the forecasts themselves are similar if two vectors are close. The models HFB24-24 and FRDM-12, for instance, share little in common when it comes to the underlying form and theoretical assumptions, though their predictive capability is very similar due to their parametric expressivity. The UNEDF series of interactions (UNEDF0, UNEDF1, UNEDF2) is another interesting case study in that they all have a very similar functional form, but are based on different calibrations. This difference directly manifests in the projections where UNEDF0 is nearly orthogonal to the other two models in the same family despite their close functional relationship. Additional information could likely be gleaned from a more targeted study of certain nuclei, but the global dataset of nuclear binding energies does not immediately reveal model specific physical insights.
@@
We now proceed to create the combined BMC+PCA model $f^{\dagger}$ as specified in Eq. (8). To this end, we analyze the RMSE performance by maximizing the likelihood (11) of the combined model as we vary the number of principal components kept. Based on the result shown in Fig. 3(c), we retain $p=7$ principal components for the combined model for the rest of this analysis. Indeed, as the number of components is increased beyond 7, the validation and test errors grow, suggesting overfitting. The dashed lines show the RMSE obtained when combining all 15 models without projecting on principal components, which also shows signs of overfitting: a lower RMSE for the training dataset and a higher RMSE for the test dataset. It is to be noted that even though the singular values shown in the inset of panel (c) decay nearly exponentially, they do not experience a rapid drop as in Fig. 2(b). In this case, the RMSE($p$) behavior for the validation set provides a good metric to determine the number of principal components that are needed to optimally model the experimental values outside of the training set.
@@
From the distributions of the weights (10), we compute the predictions of the combined model $f^{\dagger}$ with quantified uncertainties across the entire dataset. Figure 4(a) shows the predictions and $95\%$ credible intervals for the Sn ($Z=50$) isotopic chain for the unconstrained and simplex-constrained variants. One feature to note is that the simplex-constrained model has a systematically broader credible interval in its forecasts both inside and outside the training region, yet both models seem to cover the experimental data well within their credible bands. Since the constrained model can only reproduce a subset of possible combinations, we expect its performance in terms of RMSE to be less expressive in both interpolation and extrapolation, at the gain of less sensitivity to overfitting. Indeed, as can be seen in the last two rows of Table 2, the RMSE scores for the constrained model are significantly worse than for the unconstrained model - as well as the two best performing models FRDM-12 and HFB-24. Yet, they remain steady in the transition from training to validation and testing, while for the unconstrained approach the testing RMSE increases by about $20\%$ in comparison to the training RMSE.
@@
By analyzing the empirical coverage of our calibrated model, we can quantitatively assess signs of overfitting. Figure 4(b) shows the empirical coverage probability (ECP) [48, 49] for both the unconstrained and simplex-constrained models across the three datasets considered. The fact that the empirical curves all lie close to the diagonal reference line gives us confidence that the combined predictions are neither being over confident (too small credible intervals) or over conservative (too big credible intervals). This is particularly reassuring for the test set, in which considerable extrapolations have been made (cf. Fig. 3(a)). When considering just one data type (here: nuclear binding energies), the risk for overfitting is low, yet if one wishes to consider model performance on quantities not in the original dataset, the risk for overfitting can be substantially increased.
Discussion
@@
In this work, we propose, implement, and benchmark a Bayesian model combination framework accompanied by model orthogonalization using Principal Component Analysis. We discuss the features of the proposed BMC+PCA method by applying it to global models of nuclear binding energy. Following the tests based on the analytic Liquid Drop Model, we carry out realistic BMC+PCA calculations of nuclear binding energies using 15 global computational nuclear models. We demonstrate that the BMC+PCA framework performs excellently in terms of prediction accuracy and uncertainty quantification. While we have focused on the nuclear physics use case in this work, the method itself is completely general and can be applied broadly where model forecasts are utilized. It is also easy to implement and results in an interpretable combination, simplifying the application to problems that span disciplines. Furthermore, the computational scheme is robust against model repetitions and it does not favor one model class when multiple copies are added. The BMC+PCA technique can be reduced to BMM+PCA by imposing the simplex condition. In this case, only the several best performing models remain in the combination.
@@
In addition to producing optimal multimodel forecasts with robust uncertainties, BMC+PCA is also capable of identifying model collinearities. This allows the framework to perform model selection if the exact model happens to be present in the set of models. The computational efficiency of the combined model also positions it well for wide distribution on web-based platforms, such as the Bayesian Mass Explorer project [50]. While live evaluation of most of the individual models is impossible due to their inherent numerical complexity, the combined model can be evaluated, with uncertainties, on the fly. The model combination procedure itself is also efficient, meaning interested users can provide their own datasets and update the resultant BMC+PCA model.
@@ Acknowledgements
Author contributions
V.K. and W.N. conceived the project. P.G, K.G., and V.K. performed the PCA analysis. All authors discussed the results and prepared the manuscript.
Author information
Reprints and permissions information is available at www.nature.com/reprints. The authors declare no competing financial interests. Readers are welcome to comment on the online version of the paper. Correspondence and requests for materials should be addressed to W.N. (witek@frib.msu.edu)
Competing interests statement
The authors declare that they have no competing financial interests.

### v2

@@ IV.2 Case Study II: Realistic nuclear mass models
We now turn to a set of realistic models of nuclear binding energy. For this study, we have chosen 15 realistic computational models that represent a few classes of theoretical frameworks that see broad use. The models are specified in Table 2, with their respective RMSE for each of the three datasets. As the model orthogonalization and mixing strategies only require precomputed data across a range of nuclei, we pull forecasts directly from published datasets. The specific mass models chosen and their parameters are those from the MassExplorer database  [36] and HFB-24 [37] and FRDM-12  [38] mass tables.
@@
The domains of the experimental datasets used are shown in Fig. 3(a). As in the case study I, we divide the 629 data points into training, validation, and testing sets. We perform the SVD on the forecasts produced by the set of 15 models restricted to the training set and show their projections $\nu_{1}$ and $\nu_{2}$ on the first two principal components in Fig. 3(b). The nearly-exponential decay of the singular values (III) is shown in the inset of panel (c). Note that the SVD is performed after the centering procedure, thus the principal components in panel (b) should be interpreted as the variability about the average in the space of predictions. While it is tempting to draw conclusions on model similarity from these projections, it can only be said that the forecasts themselves are similar if two vectors are close. The models HFB-24 and FRDM-12, for instance, share little in common when it comes to the underlying form and theoretical assumptions, though their predictive capability is very similar due to their parametric expressivity. The UNEDF series of interactions (UNEDF0, UNEDF1, UNEDF2) is another interesting case study in that they all have a very similar functional form, but are based on different calibrations. This difference directly manifests in the projections where UNEDF0 is nearly orthogonal to the other two models in the same family despite their close functional relationship. Additional information could likely be gleaned from a more targeted study of certain nuclei, but the global dataset of nuclear binding energies does not immediately reveal model specific physical insights.
@@
We now proceed to create the combined BMC+PCA model $f^{\dagger}$ as specified in Eq. (8). To this end, we analyze the RMSE performance by maximizing the likelihood (III.1) of the combined model as we vary the number of principal components kept. Based on the result shown in Fig. 3(c), we retain $p=7$ principal components for the combined model for the rest of this analysis. Indeed, as the number of components is increased beyond 7, the validation and test errors grow, suggesting overfitting. The dashed lines show the RMSE obtained when combining all 15 models without projecting on principal components, which also shows signs of overfitting: a lower RMSE for the training dataset and a higher RMSE for the test dataset. It is to be noted that even though the singular values shown in the inset of panel (c) decay nearly exponentially, they do not experience a rapid drop as in Fig. 2(b). In this case, the RMSE($p$) behavior for the validation set provides a good metric to determine the number of principal components that are needed to optimally model the experimental values outside of the training set.
@@
From the distributions of the weights (11), we compute the predictions of the combined model $f^{\dagger}$ with quantified uncertainties across the entire dataset. Figure 4(a) shows the predictions and $95\%$ credible intervals for the Sn ($Z=50$) isotopic chain for the unconstrained and simplex-constrained variants. One feature to note is that the simplex-constrained model has a systematically broader credible interval in its forecasts both inside and outside the training region, yet both models seem to cover the experimental data well within their credible bands. Since the constrained model can only reproduce a subset of possible combinations, we expect its performance in terms of RMSE to be less expressive in both interpolation and extrapolation, at the gain of less sensitivity to overfitting. Indeed, as can be seen in the last two rows of Table 2, the RMSE scores for the constrained model are significantly worse than for the unconstrained model - as well as the two best performing models FRDM-12 and HFB-24. Yet, they remain steady in the transition from training to validation and testing, while for the unconstrained approach the testing RMSE increases by about $20\%$ in comparison to the training RMSE. That the unconstrained model $f^{\dagger}$ outperforms each individual model in terms of RMSE is not surprising – an unconstrained combination of PCs fitted to a given dataset will outperform what each individual model can do alone, as has been shown for nuclear mass models in Ref. [22]. Embedding the PCA-based model combination within a full Bayesian framework and the classification of data into training, validation, and test sets then allows for more reliable forecasts with quantified uncertainties when extrapolating beyond experimentally known masses.
@@
Indeed, by analyzing the empirical coverage of our calibrated model, we can quantitatively assess signs of overfitting. Figure 4(b) shows the empirical coverage probability (ECP)  [52, 53] for both the unconstrained and simplex-constrained models across the three datasets considered. The fact that the empirical curves all lie close to the diagonal reference line gives us confidence that the combined predictions are neither being over confident (too small credible intervals) or over conservative (too big credible intervals). This is particularly reassuring for the test set, in which considerable extrapolations have been made (cf. Fig. 3(a)). When considering just one data type (here: nuclear binding energies), the risk for overfitting is low, yet if one wishes to consider model performance on quantities not in the original dataset, the risk for overfitting can be substantially increased.
V Conclusions and Outlook
@@
In this work, we propose, implement, and benchmark a Bayesian model combination framework accompanied by model orthogonalization using Principal Component Analysis. We discuss the features of the proposed BMC+PCA method by applying it to global models of nuclear binding energy. Following the tests based on the analytic Liquid Drop Model, we carry out realistic BMC+PCA calculations of nuclear binding energies using 15 global computational nuclear models. We demonstrate that the BMC+PCA framework performs excellently in terms of prediction accuracy and uncertainty quantification. While we have focused on the nuclear physics use case in this work, the method itself is completely general and can be applied broadly where model forecasts are utilized. It is also easy to implement and results in an interpretable combination, simplifying the application to problems that span disciplines. Furthermore, the computational scheme is robust against model repetitions and it does not favor one model class when multiple copies are added. The BMC+PCA technique can be reduced to BMM+PCA by imposing the simplex condition. In this case, only the several best performing models remain in the combination, and we recover an interpretation of the model combination that can be compared to other traditional BMA and BMM approaches where the positive model weights are determined by the data. For this simplex constrained version both the Root Mean Squared Error (Table 2) as well as its uncertainty bands (Figure 4) are bigger than the unconstrained version, yet the simplex constrained approach shows signs of less over-fitting in terms of extrapolation, with a performance that remains stable across the three sets (Panel (a) of Figure 3).
@@
In addition to producing optimal multimodel forecasts with robust uncertainties, BMC+PCA is also capable of identifying model collinearities and redundancies, a functionality that stands to benefit other applications that aim to combine the wisdom of multiple nuclear models [54, 55, 56]. Furthermore, the framework is also able to perform model selection if the exact model happens to be present in the set of models. The computational efficiency of the combined model also positions it well for wide distribution on web-based platforms, such as the Bayesian Mass Explorer project [57]. While live evaluation of most of the individual models is impossible due to their inherent numerical complexity, the combined model can be evaluated, with uncertainties, on the fly. The model combination procedure itself is also efficient, meaning interested users can provide their own datasets and update the resultant BMC+PCA model.